OpenStudy (anonymous):

What is the distance between the points (2, 8) and (−6, −7)?

OpenStudy (anonymous):

\[\sqrt{(2-(-6))^2 + (8-(-7))^2} = \sqrt{8^2 + 15^2} = \sqrt{64+225}=\sqrt{289} = 17\]

OpenStudy (anonymous):

What is the distance between the points (2, 6) and (5, 2)?

OpenStudy (amistre64):

its the same process

OpenStudy (anonymous):

Yeah, just follow it.